if the hyoptenuse of a triangle measures 40, one leg is 15, what is the length of the other leg?
Answer by mananth(12270)

hypotenuse = 40
one leg = 15
leg2 ^2 = hypotenuse^2-leg1 ^2
= 40^2-15^2
=1600-225
=1375
leg2 = 37.08